28.0 g of magnesium carbonate reacts with sulfuric acid. What mass of water is produced? Show the solution.

1
Expert's answer
2022-05-30T13:29:02-0400

MgCO3 + H2SO4 = MgSO4 + H2O

According to the above reaction, 1mol (84 g) of MgCO3 reacts to form 1 mol (18 g) of H2O (water). We find how much H2O is formed when 28 g of MgCO3 reacts:

84gr MgCO3 — 18gr H2O

28gr MgCO3 — X gr H2O

X = 28 * 18 / 84 = 6 gr

ANSWER: 6 gr H2O